OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = (x – 1)^2 – 16?

OpenStudy (ybarrap):

y intercept is the value of y when x = 0 x intercept is the value of x when y = 0 y-intercept: y = (0 - 1)^2 - 16 = (-1)^2 - 16 = 1 - 16 = -15 y-intercept: -15 x-intercept y = (x – 1)^2 – 16 (x-1)^2 = y + 16 (x-1) = +/- sqrt(y + 16) x = 1 +/- sqrt(y + 6) x = 1 + sqrt(0 + 6) = 1 + sqrt(6) and x = 1 - sqrt(6) x-intercepts: 1 + 6^(1/2) and 1 - 6^(1/2) Ans. y-intercept = -15, x-intercepts = 1 + 6^(1/2) and 1 - 6^(1/2)